OpenStudy (anonymous):

When finding the sum of the arithmetic sequence 4 + 1 + –2 + –5 + –8 + –11, you will have a step of work that contains the number –3. Use complete sentences to describe how the –3 was calculated and what it means.

OpenStudy (experimentx):

think of number that contains 3 divisible by 3 13, 23 ... i guess it doesn't xx = 2(mod 3)

OpenStudy (anonymous):

to find the next term in your sequence you have to sum -3 to the previuous one

OpenStudy (experimentx):

i guess that would be 23, but ...

OpenStudy (experimentx):

man how do you calculate this algebraically??

OpenStudy (anonymous):

in your case you have 6 elements of the sequence. The sum of the 2 extream elements is always -7. 6 elements in total so Sum = -7*6/2= -21

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

back

OpenStudy (anonymous):

that's how

OpenStudy (experimentx):

so you moved the first digit of sequence by 1 ... in another 10th place huh??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

if you start with 2 estream elements (1º and last) the times you have -3 sumed there is 5. Later advance to next 2 elements that are 2º and 1 befor the last. The times -3 is sumed there is 5. and so on.....

OpenStudy (anonymous):

thnksssss

OpenStudy (anonymous):